OpenStudy (anonymous):

A man traveled to his country home, a distance of 150 miles, and then back. His average rate of speed going was 50 miles an hour and his average return speed was 30 miles an hour. His average rate of speed for the entire trip was

OpenStudy (anonymous):

I dont know if this is the right answer but.. 2(50)(30)/(50+30)=300/80=3.75mph Please help!

OpenStudy (dumbcow):

pretty close... it should be "3000" on top 3.75 mph is walking slow :)

OpenStudy (anonymous):

ohh.. i multiplied wrong. thanks

OpenStudy (dumbcow):

yw